# Saturability of the Quantum Cramér-Rao Bound in Multiparameter Quantum Estimation at the Single-Copy Level

### v1

@@ Abstract
The quantum Cramér-Rao bound (QCRB) as the ultimate lower bound for precision in quantum parameter estimation is only known to be saturable in the multiparameter setting in special cases and under conditions such as full or average commutavity of the symmetric logarithmic derivatives (SLDs) associated with the parameters. Moreover, for general mixed states, collective measurements over infinitely many identical copies of the quantum state are generally required to attain the QCRB. In the important and experimentally relevant single-copy scenario, a necessary condition for saturating the QCRB in the multiparameter setting for general mixed states is the so-called partial commutativity condition on the SLDs. However, it is not known if this condition is also sufficient. This paper derives new necessary conditions that imply partial commutativity and are almost sufficient. It is shown that together with another condition they become sufficient for saturability of the QCRB in the multiparameter single-copy case. Moreover, when the sufficient conditions are satisfied an optimal measurement saturating the QCRB can be chosen to be projective and explicitly characterized. An example is developed to illustrate the case of a multiparameter quantum state where the conditions derived herein are satisfied and can be explicitly verified.
@@ 1 Introduction
In physical systems, information about the parameters are typically obtaining by performing some measurements on a system and constructing an estimator for the parameters based on the measurement results. In quantum systems, there is an inherent fundamental noise always present, quantum noise, that persists even if all classical noise sources can be completely eliminated. Thus there has been much interest in parameter estimation when the limiting factor is quantum noise and to achieve the ultimate estimation precision physically possible, typically in the mean square sense. Quantum parameter estimation theory originated in the pioneering works of Helstrom, Holevo and Belavkin in the ’60s and ’70s and in recent years has attracted more attention as one of theoretical underpinnings for the field of quantum metrology; see [3] for a recent survey. This field aims to exploit quantum effects and quantum devices to perform measurements more accurately for emerging applications such as quantum sensing and imaging, in various physical platforms such as quantum optics, photonics and cold atoms [3, 4, 5].
@@
In the single-copy case that is of interest in the present work, a well-known result is that the QCRB can be saturated in multiparameter estimation on pure quantum states, provided that an average commutativity condition on the symmetric logarithmic derivatives (SLDs) associated with the parameters is satisfied; see Section 2 for details. For mixed states that are full rank the QCRB can be saturated if and only if the SLDs for the different parameters are mutually commuting. For general mixed states that are between these two extremes, the work [11] derived another type of commutativity condition on the SLDs, called the partial commutativity condition, and show that this condition is necessary for saturation of the QRCB in the multiparameter and single-copy case. The paper also derives necessary and sufficient conditions for general quantum measurements described by positive operator-valued measures (POVMs) to saturate the QCRB, however it does not establish the existence of such POVMs for a given quantum state. This single-copy result generalizes an analogous result in [12] for the special case of pure states and projective measurements. An alternative set of necessary and sufficient conditions for a POVM to saturate the QRRB, but expressed in terms of both the POVM and the associated estimator $\hat{\theta}$, is given in [13, Appendix B]. Whether the partial commutativity condition is sufficient has up to now been unknown. Obtaining necessary and sufficient conditions for saturating the QRCB in terms of the SLDs, if they exist, is of significance interest in practice. For instance, they can be used to determine if saturation can be achieved with fewer experimental resources.
@@
This work builds on and extends the approach of [11] to derive new necessary conditions for saturability of the QCRB in the multiparameter and single-copy setting that imply partial commutativity and are also almost sufficient (Theorem 6). They become sufficient with the addition of another condition. The proof of the theorem is constructive. When sufficient conditions for saturability are met, the proof gives an explicit construction of a measurement that saturates the QCRB, which turns out can be chosen to be projective.
@@
The paper is structured as follows. Section 2 gives a brief overview of quantum parameter estimation theory, including a statement of the QCRB and definitions of POVMs and SLDs. Section 3 reviews existing results on saturability of the QCRB in the single-copy case as well as in the more general multi-copy setting. This section also recalls some of the key results from [11] that are relevant for deriving Theorem 6. Section 4 then states and derives the main results of the paper followed by a discussion and an example of a parametrised mixed quantum state where the newly obtained conditions can be verified explicitly and full commutativity of the SLDs does not hold. Finally, Section 5 gives a summary of the content of the paper and directions for future work.
@@
Notation. $\mathbb{R}$ and $\mathbb{C}$ denote the set of real and imaginary numbers, respectively. An element of $\mathbb{R}$ or $\mathbb{C}$ is represented as a column vector unless stated otherwise. The symbol $i$ denotes the imaginary basis $i=\sqrt{-1}$. The conjugate of a complex number $c$, its real part and imaginary part will be denoted by $\overline{c}$, $\Re\{c\}$ and $\Im\{c\}$, respectively. The transpose of a matrix $X$ is denoted by $X^{\top}$ and the adjoint of an operator $X$ on a Hilbert space $\mathcal{H}$ or the conjugate transpose of a complex matrix $X$ is denoted by $X^{{\dagger}}$. A vector in a complex Hilbert space will be denoted by the ket $|x\rangle$ and its conjugate transpose by the bra $\langle x|$. The trace of a square matrix $X$ is denoted by $\mathrm{tr}(X)$. The direct sum of two vector spaces $V_{1}$ and $V_{2}$ is denoted by $V_{1}\oplus V_{2}$. For any Hermitian matrix $X$, $X\geq 0\,(>0)$ denotes that $X$ is positive semidefinite (positive definite) while $A\geq B\,(A>B)$ for any two Hermitian matrices of the same dimension denotes that $A-B\geq 0$ ($A-B>0)$. For two square matrices $X$ and $Y$, $[X,Y]=XY-YX$ and $\{X,Y\}=XY+YX$ are their commutator and anti-commutator, respectively. An $n\times n$ identity matrix will be denoted by $I_{n}$ or simply by $I$ if its dimension can be inferred from the context. Similarly, $0_{m\times n}$ will denote a zero matrix of dimension $m\times n$ with the subscript dropped if the dimension can be inferred from context. The expectation operator will be denoted by $\mathbb{E}\left[\cdot\right]$ and the expectation of a random variable $X$ by $\mathbb{E}[X]$.
@@ 2 Preliminaries
$F_{\theta,jk}=\mathrm{tr}(\rho_{\theta}\{L_{\theta_{j}},L_{\theta_{k}}\}),$
@@ 3 Overview of existing results
In this work, we are interested in the saturability of the QCRB in multiparameter quantum estimation at the single-copy level. When $\rho_{\theta}$ is full rank then the QCRB is saturable on a single copy if and only if the full commutativity condition $[L_{\theta_{j}},L_{\theta_{k}}]=0$ holds for all $j,k$. If $\rho_{\theta}$ is not full rank then the full commutativity is sufficient but no longer necessary. For pure quantum states $\rho_{\theta}=|\psi\rangle\langle\psi|$, the QCRB can be saturated on a single copy if and only the average commutativity condition holds [15, 12],
@@
$\langle\psi_{m,\theta}|[L_{\theta_{j}},L_{\theta_{k}}]|\psi_{n,\theta}\rangle,\,\forall j,k=1,\ldots,p,\,m,n=1,\ldots,r_{+}.$ (7)
@@
The partially commutativity condition (7) as a necessary condition follows from the following result [11, Theorems 1 and 2].
@@
[11, Theorem 3] If the conditions of Theorem 2 are satisfied then the partial commutativity condition (7) holds.
@@ 4 Main results and discussion
In this section, the main theorem of the paper will be stated and proven. When the conditions of theorem are satisfied, the QCRB is saturated by a projective measurement that can be explicitly characterized. An example is also developed in this section to illustrate the application of the conditions to a non-full rank quantum state with two parameters.
To get to the core arguments and methodology with minimal technicalities, the focus is on the finite-dimensional setting. However, it is reasonable to expect that the results will continue to hold, perhaps with the addition of some technical caveats, to infinite-dimensional separable Hilbert spaces when all operators have discrete countable spectra and eigenvectors. It is also reasonable to expect that the results can be extended to continuous-variable quantum systems such as Gaussian quantum systems. We begin with the following lemma.
@@
Proof. By using the representation (5) for $\rho_{\theta}$, we have that $\mathrm{tr}(\rho_{\theta}E_{k})=\mathrm{tr}(E_{k,++}\rho_{\theta,++}).$ Since $\rho_{\theta,++}>0$ and diagonal in the basis $\mathcal{B}_{+,\theta}$, this quantity is positive and $E_{k}$ is regular if and only if $E_{k,++}\geq 0$ and $E_{k,++}\neq 0$. No other conditions are imposed on $E_{k,+0}$, $E_{k,0+}=E_{k,+0}^{{\dagger}}$ and $E_{k,00}$.
On the other hand, for a null POVM operator $\mathrm{tr}(\rho_{\theta}E_{k})=0$ only if $E_{k,++}=0$. However, this condition is not sufficient for $E_{k}$ to be null since $E_{k}$ must also be positive semidefinite. Let $z$ be an $n_{s}$-dimensional complex vector, $z=[\begin{array}[]{cc}x^{\top}&y^{\top}\end{array}]$ with $x\in\mathbb{C}^{r_{+}}$ and $y\in\mathbb{C}^{r_{0}}$. Since $E_{k,++}=0$, it follows that $z^{{\dagger}}E_{k}z=2\Re\{x^{{\dagger}}E_{k,+0}y\}+y^{{\dagger}}E_{k,00}y$ and therefore $z^{{\dagger}}E_{k}z\geq 0$ for all $z$ if and only if $E_{k,+0}=0$. This proves the necessary and sufficient conditions for $E_{k}$ to be null as claimed.
Note that the conditions in Theorem 2 are actually statements about subspaces since conditions (8) and (9) hold independently of the index $n$. Indeed, it is immediately verified that (8) and (9) continue to hold when $|\psi_{n,\theta}\rangle$ is replaced by any $|\psi\rangle=\sum_{n=1}^{r_{+}}\lambda_{k}|\psi_{n,\theta}\rangle\in\mathcal{H}_{+,\theta}$ for any complex constants $\lambda_{1},\ldots,\lambda_{r+}$.
From these observations, the following statement can be extracted.
@@
The main result of this paper is Theorem 6 below. The main idea of the proof is to show that the conditions stated in the theorem are necessary to satisfy the conditions of Theorem 2 and are sufficient under an additional condition. The proof involves the construction of an optimal POVM.
@@
The QCRB is saturable at the single-copy level only if the following conditions hold:
  1. 1.
@@
  2. 2.
    $L_{\theta_{m},+0}L_{\theta_{l},+0}^{{\dagger}}-L_{\theta_{l},+0}L_{\theta_{m},+0}^{{\dagger}}=0$ for all $l,m=1,\ldots,p$.
@@
Conversely, if Conditions 1 and 2 are satisfied and, in addition, if all corresponding columns of $L_{\theta_{l},+0}$ and $L_{\theta_{m},+0}$ with $l,m=1,\ldots,p$ are real scalar multiples of one another or the corresponding columns are simultaneously vanishing, that is, the $s$-th column of $L_{\theta_{l},+0}$ is either $\lambda_{lms}\in\mathbb{R}$ times the $s$-th column of $L_{\theta_{m},+0}$, or both columns are zero, for all columns $s$ and $\forall l,m=1,\ldots,p$ then the QCRB is saturated. When these conditions are satisfied, there exists an optimal projective measurement given by the POVM:
@@
$\underbrace{\biggl{\{}\left[\begin{array}[]{cc}\Pi_{\theta,1}&0\\ 0&0\end{array}\right],\ldots,\left[\begin{array}[]{cc}\Pi_{\theta,\chi_{\theta}}&0\\ 0&0\end{array}\right]\biggr{\}}}_{\hbox{Regular POVM operators}}\bigcup\underbrace{\biggl{\{}\left[\begin{array}[]{cc}0&0\\ 0&E_{1,00}\end{array}\right],\ldots,\left[\begin{array}[]{cc}0&0\\ 0&E_{r_{0},00}\end{array}\right]\biggr{\}}}_{\hbox{Null POVM operators}},$
@@
where $\Pi_{\theta_{j}}$ for $j=1,\ldots,\chi_{\theta}$ ($\chi_{\theta}\leq r_{+}$) are the (common) projection operators in the spectral decomposition of $L_{\theta_{l},++}$ for $l=1,\ldots,p$, and $E_{j,00}$ is the $r_{0}\times r_{0}$ projection operator which is zero everywhere except for a 1 in row $j$ and column $j$ for $j=1,\ldots,r_{0}$.
To prove the theorem, the following lemma will be used.
Lemma 7
If $A$ and $B$ are two complex matrices of the same dimension then $AB^{{\dagger}}-B^{{\dagger}}A=0$ if all corresponding non-zero columns of $A$ and $B$ are real scalar multiples of one another.
Proof. Let $X_{j}$ denote the $j$-th column of $X$ ($X$ is either $A$ or $B$) and let $X_{jk}$ be the $k$-th element of $X_{j}$. Denote the number of rows by $R$ and the number of columns by $C$. We have that $(AB^{{\dagger}}-B^{{\dagger}}A)\alpha=0$ for all $\alpha\in\mathbb{C}^{R}$, where $\alpha=(\alpha_{1},\ldots,\alpha_{R})^{\top}$. This is equivalent to
$\sum_{j=1}^{R}\alpha_{j}\sum_{k=1}^{C}(A_{k}\bar{B}_{kj}-B_{k}\bar{A}_{kj})=0,$
for all $\alpha_{1},\ldots,\alpha_{R}$. It can be rewritten as
$\sum_{k=1}^{C}\left(A_{k}\left(\sum_{j=1}^{R}\alpha_{j}\bar{B}_{kj}\right)-B_{k}\left(\sum_{j=1}^{R}\alpha_{j}\bar{A}_{kj}\right)\right)=0.$ (12)
If one of $A_{k}$ or $B_{k}$ is zero then they do not contribute to the sum on the left hand side of (12) and can both be ignored. Therefore for $(AB^{{\dagger}}-B^{{\dagger}}A)\alpha=0$ $\forall\alpha\in\mathbb{C}^{R}$, it is enough that $\sum_{j=1}^{R}\alpha_{j}\bar{B}_{kj}=0$ and $\sum_{j=1}^{R}\alpha_{j}\bar{A}_{kj}=0$ simultaneously for all $\alpha_{1},\ldots,\alpha_{R}$ and all $k$ for which $A_{k}$ and $B_{k}$ are non-vanishing. In particular, it is sufficient that $B_{k}=c_{k}A_{k}$ for some constant $c_{k}$ for all non-zero $A_{k}$ and $B_{k}$. Substituting this into (12),
$02i\ \sum_{k=1}^{C}\Im\{c_{k}\}A_{k}\left(\sum_{j=1}^{R}\alpha_{j}\bar{A}_{kj}\right)=$ 0 .
Therefore, for each non-vanishing column pairs $A_{k}$ and $B_{k}$ the associated non-zero constant $c_{k}$ must be real for the sum to vanish as required.
We can now proceed with a proof of Theorem 6.
@@
Proof of Theorem 6. We begin by noting that the partial commutativity condition (7) can be equivalently stated as
@@
Using the four block decomposition (2) of $L_{\theta_{j}}$ as an observable, a simple calculation shows that the above identity is equivalent to:
$[L_{\theta_{l},++},L_{\theta_{m},++}]+(L_{\theta_{l},+0}L_{\theta_{m},+0}^{{\dagger}}-L_{\theta_{l},+0}L_{\theta_{m},+0}^{{\dagger}})=0.$ (13)

### v3

@@ Abstract
The quantum Cramér-Rao bound (QCRB) as the ultimate lower bound for precision in quantum parameter estimation is only known to be saturable in the multiparameter setting in special cases and under conditions such as full or average commutavity of the symmetric logarithmic derivatives (SLDs) associated with the parameters. Moreover, for general mixed states, collective measurements over infinitely many identical copies of the quantum state are generally required to attain the QCRB. In the important and experimentally relevant single-copy scenario, a necessary condition for saturating the QCRB in the multiparameter setting for general mixed states is the so-called partial commutativity condition on the SLDs. However, it is not known if this condition is also sufficient. This paper establishes necessary and sufficient conditions for saturability of the multiparameter QCRB in the single-copy setting in terms of the commutativity of a set of projected SLDs and the existence of a unitary solution to a system of nonlinear partial differential equations. New necessary conditions that imply partial commutativity are also obtained, which together with another condition become sufficient. Moreover, when the sufficient conditions are satisfied an optimal measurement saturating the QCRB can be chosen to be projective and explicitly characterized. An example is developed to illustrate the case of a multiparameter quantum state where the conditions derived herein are satisfied and can be explicitly verified. [This work will be published in IEEE Control Systems Letters (DOI: 10.1109/LCSYS.2024.3382330)]
@@ 1 Introduction
In physical systems, information about the parameters are typically obtaining by performing some measurements on a system and constructing an estimator for the parameters based on the measurement results. In quantum systems, there is an inherent fundamental noise always present, quantum noise, that persists even if all classical noise sources can be completely eliminated. Thus there has been much interest in parameter estimation when the limiting factor is quantum noise and to achieve the ultimate estimation precision physically possible, typically in the mean square sense. Quantum parameter estimation theory originated in the pioneering works of Helstrom, Holevo and Belavkin in the ’60s and ’70s and in recent years has attracted more attention as one of theoretical underpinnings for the field of quantum metrology; see [3] for a recent survey. This field aims to exploit quantum effects and quantum devices to perform measurements more accurately for emerging applications such as quantum sensing and imaging, in various physical platforms such as quantum optics, photonics and cold atoms [3, 4, 5]. In quantum systems and control, quantum parameter estimation is foundational for quantum system identification; see, e.g., [6] and the references therein.
@@
In the single-copy case that is of interest in the present work, a well-known result is that the QCRB can be saturated in multiparameter estimation on pure quantum states, provided that an average commutativity condition on the symmetric logarithmic derivatives (SLDs) associated with the parameters is satisfied; see Section 2 for details. For mixed states that are full rank the QCRB can be saturated if and only if the SLDs for the different parameters are mutually commuting [12, ref. [5] and Appendix B.1]. For general mixed states that are between these two extremes, the work [13] derived another type of commutativity condition on the SLDs, called the partial commutativity condition, and show that this condition is necessary for saturation of the QRCB in the multiparameter and single-copy case. The paper also derives necessary and sufficient conditions for general quantum measurements described by positive operator-valued measures (POVMs) to saturate the QCRB, however it does not establish the existence of such POVMs for a given quantum state. This single-copy result generalizes an analogous result in [14] for the special case of pure states and projective measurements. An alternative set of necessary and sufficient conditions for a POVM to saturate the QRRB, but expressed in terms of both the POVM and the associated estimator $\hat{\theta}$, is given in [12, Appendix B.1]. Whether the partial commutativity condition is sufficient has up to now been unknown. Obtaining necessary and sufficient conditions for saturating the QRCB in terms of the SLDs, if they exist, is of significance interest in practice. For instance, they can be used to determine if saturation can be achieved with fewer experimental resources.
@@
This paper establishes necessary and sufficient conditions for saturability of the multiparameter QCRB in the single-copy setting. Then new necessary conditions that imply partial commutativity are obtained, building on the approach of [13]. They become sufficient with the addition of another condition (Theorem 6). When sufficient conditions for saturability are met, the proof gives an explicit construction of a measurement that saturates the QCRB, which turns out can be chosen to be projective.
@@
The paper is structured as follows. Section 2 gives a brief overview of quantum parameter estimation theory, including a statement of the QCRB and definitions of POVMs and SLDs. Section 3 reviews existing results on saturability of the QCRB in the single-copy case as well as in the multi-copy setting. This section also recalls some of the key results from [13] that are relevant for deriving Theorem 6. Section 4 then states and derives the main results of the paper followed by a discussion and an example. Finally, Section 5 gives a conclusion for the paper and directions for future work.
@@
Notation. $\mathbb{R}^{n}$ and $\mathbb{C}^{n}$ denote the set of real and complex vectors, respectively, whose elements are represented as a column vector unless stated otherwise, and $i=\sqrt{-1}$. The conjugate of a complex number $c$, its real part and imaginary part will be denoted by $\overline{c}$, $\Re\{c\}$ and $\Im\{c\}$, respectively. The transpose of a matrix $X$ is denoted by $X^{\top}$ and the adjoint of an operator $X$ on a Hilbert space $\mathcal{H}$ or the conjugate transpose of a complex matrix $X$ is denoted by $X^{{\dagger}}$. A vector in a complex Hilbert space will be denoted by the ket $|x\rangle$ and its conjugate transpose by the bra $\langle x|$. The trace of a square matrix $X$ is denoted by $\mathrm{tr}(X)$. The direct sum of two vector spaces $V_{1}$ and $V_{2}$ is denoted by $V_{1}\oplus V_{2}$. For a vector space $V$ and an operator $X$ on a vector space containing $V$, $\left.X\right|_{V}$ denotes the restriction of $X$ to $V$. For any Hermitian matrix $X$, $X\geq 0\,(>0)$ denotes that $X$ is positive semidefinite (positive definite) while $A\geq B\,(A>B)$ for any two Hermitian matrices of the same dimension denotes that $A-B\geq 0$ ($A-B>0)$. For two square matrices $X$ and $Y$, $[X,Y]=XY-YX$ and $\{X,Y\}=XY+YX$ are their commutator and anti-commutator, respectively. An $n\times n$ identity matrix will be denoted by $I_{n}$ or simply by $I$ if its dimension can be inferred from the context. Similarly, $0_{m\times n}$ will denote a zero matrix of dimension $m\times n$ with the subscript dropped if the dimension can be inferred from context. The expectation operator will be denoted by $\mathbb{E}\left[\cdot\right]$ and the expectation of a random variable $X$ by $\mathbb{E}[X]$.
@@ 2 Preliminaries
$F_{\theta,jk}=\mathrm{tr}(\rho_{\theta}\{L_{\theta_{j}},L_{\theta_{k}}\})/2,$
@@ 3 Overview of existing results
In this work, we are interested in the saturability of the QCRB in multiparameter quantum estimation at the single-copy level. When $\rho_{\theta}$ is full rank then the QCRB is saturable on a single copy if and only if the full commutativity condition $[L_{\theta_{j}},L_{\theta_{k}}]=0$ holds for all $j,k$. If $\rho_{\theta}$ is not full rank then the full commutativity is no longer necessary. For pure quantum states $\rho_{\theta}=|\psi\rangle\langle\psi|$, the QCRB can be saturated on a single copy if and only the average commutativity condition holds [16, 14],
@@
$\langle\psi_{m,\theta}|[L_{\theta_{j}},L_{\theta_{k}}]|\psi_{n,\theta}\rangle=0,\,\forall j,k=1,\ldots,p,\,m,n=1,\ldots,r_{+}.$ (7)
@@
The partially commutativity condition (7) as a necessary condition follows from the following result [13, Theorems 1 and 2].
@@
[13, Theorem 3] If the conditions of Theorem 2 are satisfied then the partial commutativity condition (7) holds.
@@ 4 Main results and discussion
In this section, the main theorem of the paper will be stated and proven. When sufficient conditions of the theorem are satisfied, the QCRB is saturated by a projective measurement that can be explicitly characterized. An example is also developed to illustrate the application of the theorem to a non-full rank quantum state with two parameters.
To get to the core arguments and methodology with minimal technicalities, the focus is on the finite-dimensional setting. However, it is reasonable to expect that the results will continue to hold, perhaps with the addition of some technical caveats, to infinite-dimensional separable Hilbert spaces when all operators have discrete countable spectra and eigenvectors (e.g., [13] allows an infinite-dimensional Hilbert space with a countable basis). It is also reasonable to expect that the results can be extended to continuous-variable quantum systems such as Gaussian quantum systems. We begin with the following lemma.
@@
Proof. By using the representation (5) for $\rho_{\theta}$, we have that $\mathrm{tr}(\rho_{\theta}E_{k})=\mathrm{tr}(E_{k,++}\rho_{\theta,++}).$ Since $\rho_{\theta,++}>0$ and diagonal in the basis $\mathcal{B}_{+,\theta}$, this quantity is positive and $E_{k}$ is regular if and only if $E_{k,++}\geq 0$ and $E_{k,++}\neq 0$. No other conditions are imposed on $E_{k,+0}$, $E_{k,0+}=E_{k,+0}^{{\dagger}}$ and $E_{k,00}$ other than that they are such that $E_{k}\geq 0$.
On the other hand, for a null POVM operator $\mathrm{tr}(\rho_{\theta}E_{k})=0$ only if $E_{k,++}=0$. However, this is not sufficient since $E_{k}$ must also be positive semidefinite. Let $z$ be an $n_{s}$-dimensional complex vector, $z=[\begin{array}[]{cc}x^{\top}&y^{\top}\end{array}]$ with $x\in\mathbb{C}^{r_{+}}$ and $y\in\mathbb{C}^{r_{0}}$. Since $E_{k,++}=0$, it follows that $z^{{\dagger}}E_{k}z=2\Re\{x^{{\dagger}}E_{k,+0}y\}+y^{{\dagger}}E_{k,00}y$ and therefore $z^{{\dagger}}E_{k}z\geq 0$ for all $z$ if and only if $E_{k,+0}=0$. This proves the necessary and sufficient conditions for $E_{k}$ to be null.
Note that the conditions in Theorem 2 are actually statements about subspaces since conditions (8) and (9) hold independently of the index $n$. Indeed, it is immediately verified that (8) and (9) continue to hold when $|\psi_{n,\theta}\rangle$ is replaced by any $|\psi\rangle=\sum_{n=1}^{r_{+}}\lambda_{k}|\psi_{n,\theta}\rangle\in\mathcal{H}_{+,\theta}$ for any complex constants $\lambda_{1},\ldots,\lambda_{r+}$. From these observations, the following statement can be extracted.
@@
The main result of this paper is Theorem 6 below.
@@
Consider the following four conditions:
  1. 1)
@@
  2. 2)
    For each $\theta$ there exists a unitary $U_{\theta}\in\mathbb{C}^{r_{+}\times r_{+}}$ such that $U_{\theta}^{{\dagger}}(\partial_{l}U_{\theta}-U_{\theta}V_{\theta}^{{\dagger}}\partial_{l}V_{\theta})\rho_{\theta,++}+\rho_{\theta,++}(\partial_{l}U_{\theta}-U_{\theta}V_{\theta}^{{\dagger}}\partial_{l}V_{\theta})^{{\dagger}}U_{\theta}=0$ for $l=1,\ldots,p$, where $\partial_{l}=\partial/\partial\theta_{l}$, $V_{\theta}=[\begin{array}[]{ccc}|\psi_{1,\theta}\rangle&\ldots&|\psi_{r_{+},\theta}\rangle\end{array}]$, and $\rho_{\theta,++}$ is represented in the basis $\mathcal{B}_{+,\theta}$.
  3. 3)
    $L_{\theta_{l},+0}L_{\theta_{m},+0}^{{\dagger}}-L_{\theta_{m},+0}L_{\theta_{l},+0}^{{\dagger}}=0$ for all $l,m=1,\ldots,p$.
  4. 4)
@@
    There exists an $r_{0}\times r_{0}$ complex unitary matrix $W$ such that all corresponding columns of $L_{\theta_{l},+0}W$ and $L_{\theta_{m},+0}W$ with $l,m=1,\ldots,p$ are real scalar multiples of one another or the corresponding columns are simultaneously vanishing. That is, the $s$-th column of $L_{\theta_{l},+0}W$ is either $\lambda_{lms}\in\mathbb{R}$ times the $s$-th column of $L_{\theta_{m},+0}W$, or both columns are zero (i.e., all their elements are zero), for all columns $s$ and $\forall l,m$.
Then the following statements hold for saturability of the multiparameter QCRB in the single-copy setting:
  1. i)
    Conditions 1 and 2 are necessary and sufficient.
  2. ii)
    Condition 3 is necessary and together with Condition 1 imply partial commutativity.
  3. iii)
    Conditions 1 and 4 are sufficient.
When the sufficient conditions (Conditions 1 and 4) are satisfied, there exists an optimal projective measurement given by the POVM:
@@
$\underbrace{\biggl{\{}\left[\begin{array}[]{cc}\Pi_{\theta,1}&0\\ 0&0\end{array}\right],\ldots,\left[\begin{array}[]{cc}\Pi_{\theta,\chi_{\theta}}&0\\ 0&0\end{array}\right]\biggr{\}}}_{\hbox{Regular POVM operators}}\bigcup\underbrace{\biggl{\{}\left[\begin{array}[]{cc}0&0\\ 0&WD_{00,1}W^{{\dagger}}\end{array}\right],\ldots,\left[\begin{array}[]{cc}0&0\\ 0&WD_{00,r_{0}}W^{{\dagger}}\end{array}\right]\biggr{\}}}_{\hbox{Null POVM operators}},$
@@
where $\Pi_{\theta_{j}}$ for $j=1,\ldots,\chi_{\theta}$ ($\chi_{\theta}\leq r_{+}$) are the (common) projection operators in the spectral decomposition of $L_{\theta_{l},++}$ for $l=1,\ldots,p$, and $D_{00,j}$ is the $r_{0}\times r_{0}$ projection operator which is zero everywhere except for a 1 in row $j$ and column $j$ for $j=1,\ldots,r_{0}$.
Remark 7
Note that it can be verified by direct calculation that Condition 4 of the theorem on the existence of $W$ and on the columns of $L_{\theta_{l},+0}W$ and $L_{\theta_{m},+0}W$ implies the necessary Condition 3 of the theorem. Therefore, when Conditions 1 and 4 are satisfied, Condition 3 is redundant.
@@
Proof of Theorem 6. We begin by noting that the partial commutativity condition (7) can be stated as
@@
Using the four block decomposition (2) of $L_{\theta_{j}}$ as an observable, the identity above is equivalent to:
$[L_{\theta_{l},++},L_{\theta_{m},++}]+(L_{\theta_{l},+0}L_{\theta_{m},+0}^{{\dagger}}-L_{\theta_{m},+0}L_{\theta_{l},+0}^{{\dagger}})=0.$ (12)